# AffordanceSAM: Segment Anything Once More in Affordance Grounding

### v1

@@ Appendix 6 More Details about Ablation Setup
In this section, we have a detailed illustration of the evaluation setup of each table in our ablation study. The table index here refers to the index in the main paper.
Table 3. This table presents the effect of our proposed coarse-to-fine training recipe. The first line of results is obtained by directly using EVF-SAM’s checkpoint [^3] https://github.com/hustvl/EVF-SAM to evaluate. Then we gradually add the training data divided by stage and use the last checkpoint to evaluate. When combining all the data, we add affordance-adaption module at the beginning of the training, which is a sightly different from our default setting.
Table 4. This table presents the effect of each component in our proposed affordance-adaption module. The learnable queries in the second row do not conduct any cross-attention with mentioned features but conduct self-attention twice as we mentioned in Section 3.2 in the main paper.
Table 5 and Table 6. These two tables present the effect of design choise of our proposed post-processing program. In first row of Table 5, we do not use the 5$\sim$6 steps in Algorithm 1. And so on for the second and third rows. Noticing that we only vary five values of $\gamma$ results in Table 6, we believe more optional value of $\gamma$ can be further searched; however, given the constraints of time and computational resources, as well as the diminishing returns on additional searches, we have decided not to pursue further searches. This decision does not affect the significance of our contribution.
@@ Appendix 7 Failure Cases and Feature Work
Figure 6: Failure cases when facing multiple objects or multiple affordance functions.
Although our AffordanceSAM achieves state-of-the-art (SOTA) performance under the AGD20K benchmark, and generalized well beyond the training data, when faced with more complex scenarios (e.g. muti-objects affordance), it sometimes cannot get accurate results. As shown in Figure 6, first, we find AffordanceSAM fails on completely separating two different functional regions of an object when prompted with muti-actions (see the first row of Figure 6). Next, we find that AffordanceSAM can not get an accurate affordance map when two different objects in a picture but prompted with the same affordance action (see the second row of Figure 6). These limitations may be caused by the fact that every image we use to train AffordanceSAM is single object format with single affordance region involved.
Thus, endowing AffordanceSAM with the capability to deal with more complex scenarios like multiple objects or multiple affordance functions within a single image might be an exciting avenue for future research and development.

### v2

@@ Appendix 6 More Details about Ablation Setup
In this section, we have a detailed illustration of the evaluation setup of each table in our ablation study. The table index here refers to the index in the ablation part of the main paper.
Table 3. This table presents the effect of our proposed C2F-Aff dataset and training recipe. The first line of results is obtained by directly using EVF-SAM’s checkpoint to evaluate. Then we gradually add the training data, divided by the dataset part, and use the last checkpoint to evaluate. When combining all the data, we add the affordance-adaption module at the beginning of the training, which is slightly different from our default setting.
Table 4. This table presents the effect of each component in our proposed affordance-adaption module. The learnable queries in the second row do not conduct any cross-attention with the mentioned features but conduct self-attention twice, as we mentioned in Section 3.2 in the main paper.
[…]
@@ Appendix 7 Failure Cases and Feature Work
Figure 7: Failure cases when facing multiple objects or multiple affordance actions.
Although our AffordanceSAM achieves state-of-the-art (SOTA) performance under the AGD20K benchmark, and generalizes well beyond the training data, when faced with more complex scenarios (e.g., multi-object affordance), it sometimes cannot get accurate results. As shown in Figure 7, first, we find AffordanceSAM fails on completely separating two different functional regions of an object when prompted with multi-actions (see the first row of Figure 7). Next, we find that AffordanceSAM can not get an accurate affordance map when two different objects in a picture but are prompted with the same affordance action (see the second row of Figure 7). These limitations may be caused by the fact that every image we use to train AffordanceSAM is a single object format with a single affordance region involved.
Thus, endowing AffordanceSAM with the capability to deal with more complex scenarios like multiple objects or multiple affordance actions within a single image might be an exciting avenue for future research and development.